OpenStudy (anonymous):

Can someone please help me!! ill fan you and give you a metal!! Functional Values

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (misty1212):

hi

OpenStudy (misty1212):

the reason no one answers is because there is no question to answer do you have a question? maybe i can help

OpenStudy (anonymous):

Given that (8,-1) is on the graph of f(x), find the corresponding point for the function f(-2/3 x).

OpenStudy (misty1212):

\[f(-\frac{2}{3}x)\]?

OpenStudy (anonymous):

yes that lol

OpenStudy (misty1212):

lol ok this is very easy but confusing what you need to do is solve \[\frac{-2}{3}x=8\] for \(x\)

OpenStudy (anonymous):

x=-12?

OpenStudy (misty1212):

yes, now we can answer it

OpenStudy (misty1212):

the answer is \((-12,1)\) is on the graph of \(f(-\frac{2}{3}x)\)

OpenStudy (misty1212):

because if \(x=-12\) then \[f(-\frac{2}{3}\times (-12))=f(8)=1\]

OpenStudy (anonymous):

im lost????????

OpenStudy (misty1212):

ok lets go a bit slow then

OpenStudy (misty1212):

because once you understand what the question is asking, it if very simple to solve it you solved it already quickly, by solving \(-\frac{2}{3}x=8\)

OpenStudy (misty1212):

so since you solved it, and got \(x=-12\) what does that mean? it means that \[-\frac{2}{3}\times (-12)=8\] right?

OpenStudy (anonymous):

ya

OpenStudy (misty1212):

ok and the question is asking "if \((8,1)\) is on the graph blah blah" so all you really know about the function is \(f(8)=1\) right?

OpenStudy (misty1212):

now is when you say "ya, \(f(8)=1\) and that is all we know about \(f\)" that is clear or not?

OpenStudy (anonymous):

i kinda understand the graphing stuff is just not my thing lol its way to conffusing for me /

OpenStudy (misty1212):

ok it is not really about graphing though \((8,1)\) is on the graph is another way of saying that \(f(8)=1\)

OpenStudy (misty1212):

that is how you graph, input in the first coordinate, output in the second so if \((a,b)\) is on the graph that just means \(f(a)=b\)

OpenStudy (anonymous):

okay so we know x=-12

OpenStudy (misty1212):

so what we know exactly about \(f\) is \(f(8)=1\) nothing else we are asked about \(f(-\frac{2}{3}x)\) so the only thing we can do is figure out what to plug in for \(x\) to get \(8\) that is the same as solving \[-\frac{2}{3}x=8\] for \(x\)

OpenStudy (anonymous):

and x=-12 right?

OpenStudy (misty1212):

once you find that solving \[-\frac{2}{3}x=8\] you get \[x=-12\] that means \[f(-\frac{2}{3}]\times (-12))=f(8)=1\]

OpenStudy (misty1212):

so if \((8,1)\) is on the graph of \(f(x)\) then \((-12,1)\) is on the graph of \(f(-\frac{2}{3}x)\)

OpenStudy (anonymous):

actually its 8,-1

OpenStudy (misty1212):

the simplest way to think about it if you see another one is this say \((3,5)\) is on the graph of \(f(x)\) and you are asked what is on the graph of \(f(2x+7)\) then all you wold do is solve \[2x+7=3\] and put that in the first spot

OpenStudy (misty1212):

ok then the answer is \((-12,1)\)

OpenStudy (misty1212):

what i am trying to say is that though it may be confusing, finding the solution is easy

OpenStudy (anonymous):

no it said it was wrong

OpenStudy (princeharryyy):

try (12,-1) this will be right

OpenStudy (anonymous):



OpenStudy (anonymous):

hmm i would go with @misty1212 answer of \((-12,-1)\)

OpenStudy (princeharryyy):

yup sorry! it would be (-12,-1) a bit of wrong calculation... thats right.

OpenStudy (anonymous):

she got it right just forgot the negative

OpenStudy (princeharryyy):

yup!